# Simon Briend

### v2

@@ 1 Introduction
$\widehat{\sigma}(T,\sigma)\overset{\mathcal{L}}{=}\widehat{\sigma}(T^{\sigma^{\prime}},\sigma\circ\sigma^{\prime})~{},$ (1)
@@
$R_{\alpha}(\widehat{\sigma})\ \stackrel{{\scriptstyle\rm def}}{{=}}\ \mathbb{E}\left[\sum_{i=1}^{n}\frac{|\widehat{\sigma}(i)-\sigma(i)|}{\sigma(i)^{\alpha}}\right]~{},$ (2)
@@
While the risk defined in (2) can be computed for any value $\alpha>0$, we restrict $\alpha$ to a range of values which are relevant for our ordering problem. Specifically, we only consider $\alpha\geq 1$, since for $\alpha<1$ the problem becomes trivial, since even a random permutation has a risk which is minimax optimal up to constant factor (see Appendix A.1). In Theorems 1 and 7, we provide minimax lower bounds for the risk $R_{\alpha}(\widehat{\sigma})$ in the urrt and pa model, for any label-invariant estimator $\widehat{\sigma}$. Then, in Theorems 4 and 8 upper bounds for the risk of the Jordan ordering are obtained. Finally, in Corollaries 5 and 9 we prove that our proposed estimator is minimax optimal up to constant factors, in a non-trivial range of parameters $\alpha$. In the following table, we summarise our findings. For $\alpha\geq 1$, we denote by $R^{*}_{\alpha}$ the optimal risk, and $R_{\alpha}(\widehat{\sigma}_{J})$ the risk of the Jordan ordering.
@@
In the rest of this section, we review previous works and introduce some notation. Then, in Section 2, we analyze the Jordan ordering in the urrt model. Next, we consider the pa model and prove analogous results in Section 3. Finally, in Section 4, we report the results of our simulation study and compare the empirical performance of the Jordan estimator to alternative methods based on the degree centrality, a peeling method (Navlakha and Kingsford [24, Section 2.3]) and a spectral method commonly used in seriation problems (Recanati et al. [26]).
@@ 1.1 Related work
The vertex arrival-time estimation problem bears some similarity to the seriation problem, though in the former, the dependence is intrinsically related to the tree structure. For example, in a random geometric graph (Gilbert [14]), the seriation problem is to estimate the position of the random points. Since there is no time structure in seriation, different metrics for the error are used, such as the maximum distance between the true and estimated latent position. Examples of methods are provided by Giraud et al. [15]. Another widely studied seriation method consists in ordering latent points by a spectral method on the graph Laplacian (see Section 4 and Recanati et al. [26] for details). They give guarantees for the quality of their method when the observed adjacency matrix is a perturbed Robinson matrix. The expected adjacency matrices of both urrt and pa trees are Robinson, and therefore the models studied here can be viewed as perturbed Robinson matrices. Nonetheless, none of the above-mentioned papers gives any insight about seriation in urrt and pa trees.
@@ 1.2 Notation
Recall that for a tree $T$ and a vertex $u\in T$, we denote by $(T,u)$ the tree rooted at $u$. For a rooted tree $(T,u)$ and a vertex $v$ we denote by $(T,u)_{v}$ the subtree of $T$ consisting of all vertices $w$ such that $v$ lies on the path connecting $w$ to $u$ (see Figure 1). For two vertices $u,v\in T$, $u\sim v$ means that $u$ is a neighbor of $v$ in $T$ (and reciprocally). In a rooted tree $(T,u)$, we say that $w$ is a child of $v$ if $w$ is in $(T,u)_{v}$. We denote by ${\rm de}_{n}(u)=|(\mathcal{T}_{n},1)_{u}|$ the number of descendants of $u$ in $\mathcal{T}_{n}$. For simplicity, we drop the subscript $n$ and use ${\rm de}(u)$ when the size of the tree is fixed and clear from the context.
@@ 2.1 A lower bound
$R^{*}_{\alpha}:=\min_{\widehat{\sigma}\in\Pi_{n}}R_{\alpha}(\widehat{\sigma})~{},$
@@
In the urrt model, we have, for all $\alpha>0$ and $n\geq 200$, $R^{*}_{\alpha}\geq\frac{n^{2-\alpha}}{70}~{}.$
@@
$R_{\alpha}(\widehat{\sigma})=\mathbb{E}\left[\sum_{j=1}^{n}\frac{\left|\widehat{\sigma}(j)-\sigma(j)\right|}{\sigma(j)^{\alpha}}\right]=\mathbb{E}\left[\sum_{j=1}^{n}\frac{\left|\widehat{\sigma}\circ\tau(j)-j\right|}{j^{\alpha}}\right]~{},$
@@
$\geq\frac{1}{n^{\alpha}}\sum_{j=\left\lfloor n/2\right\rfloor+1}^{\left\lfloor 3n/4\right\rfloor}\mathbb{E}\left[\big{|}\widehat{\sigma}\circ\tau(j)-j\big{|}+\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\ \left\lfloor\frac{n}{4}\right\rfloor-j\bigg{|}\right]~{}.$ (6)
@@
$\Omega_{j}:=\left\{\tau(j)\text{ and }\tau(\left\lfloor n/4\right\rfloor+j)\text{ are leaves, connected to vertices of rank }\leq n/2\right\}~{}.$
@@
    $\prod_{k=j+1}^{\left\lfloor n/4\right\rfloor+j-1}\frac{k-2}{k-1}~{}.$
@@
    $\prod_{k=\left\lfloor n/4\right\rfloor+j+1}^{n}\frac{k-3}{k-1}~{}.$
@@
$\mathbb{P}\left\{\Omega_{j}\right\}$ $=\frac{\left\lfloor n/2\right\rfloor}{j-1}\cdot\frac{\left\lfloor n/2\right\rfloor}{\left\lfloor n/4\right\rfloor+j}\cdot\prod_{k=j+1}^{\left\lfloor n/4\right\rfloor+j-1}\frac{k-2}{k-1}\cdot\prod_{k=\left\lfloor n/4\right\rfloor+j+1}^{n}\frac{k-3}{k-1}$
$=\frac{\left\lfloor n/2\right\rfloor}{j-1}\cdot\frac{\left\lfloor n/2\right\rfloor}{\left\lfloor n/4\right\rfloor+j}\cdot\frac{j-1}{\left\lfloor n/4\right\rfloor+j-2}\cdot\frac{(\left\lfloor n/4\right\rfloor+j-2)(\left\lfloor n/4\right\rfloor+j-1)}{(n-2)(n-1)}~{},$
@@
$\mathbb{P}\left\{\Omega_{j}\right\}\geq\frac{1}{4}\left(1-\frac{1}{n-2}\right)~{}.$ (7)
@@
$\geq$ $\frac{1}{4}\left(1-\frac{1}{n-2}\right)^{3}\mathbb{E}\left[\left.\big{|}\widehat{\sigma}\circ\tau(j)-j\big{|}+\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)\bigg{|}\ \right|\ \Omega_{j}\right]~{}.$
@@
$\ \ \ \ +\sum_{t\in\mathbb{T}}\mathbb{P}\left\{T=t^{\gamma}\mid\Omega_{j}\right\}\mathbb{E}\left[\left.\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)\bigg{|}\ \right|\ \Omega_{j},\ T=t^{\gamma}\right]~{},$
@@
$\mathbb{P}\left\{T=t\mid\Omega_{j}\right\}=\mathbb{P}\left\{T=t^{\gamma}\mid\Omega_{j}\right\}~{}.$
@@
$\widehat{\sigma}[T^{\gamma}]\circ\gamma\overset{\mathcal{L}}{=}\widehat{\sigma}[T]~{},$
@@
$\left(\widehat{\sigma}(j)\ \mid\Omega_{j},T=t\right)\overset{\mathcal{L}}{=}\left(\widehat{\sigma}\left(\left\lfloor n/4\right\rfloor+j\right)\ \mid\Omega_{j},T=t^{\gamma}\right)~{},$
@@
$\mathbb{E}\left[|\widehat{\sigma}\circ\tau(j)-j|\ \mid\ \Omega_{j},\ T=t\right]+\mathbb{E}\left[\left.\left|\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left\lfloor\frac{n}{4}\right\rfloor-j\right|\ \right|\ \Omega_{j},\ T=t^{\gamma}\right]\geq\lfloor n/4\rfloor~{}.$
@@
$\mathbb{E}\left[|\widehat{\sigma}\circ\tau(j)-j|+\left|\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left\lfloor\frac{n}{4}\right\rfloor-j\right|\right]\geq\frac{n-4}{16}\left(1-\frac{1}{n-2}\right)~{}.$
@@
$R_{\alpha}(\widehat{\sigma})\geq\frac{n^{2-\alpha}}{70}~{},$
@@ 2.2 An auxiliary “descendant-ordering” procedure
    $\widehat{\sigma}_{J}(1)\leq 2.5\frac{\log(1/\epsilon)}{\epsilon}~{}.$
@@
$R_{\alpha}(\widehat{\sigma}_{J})\leq R_{\alpha}(\widehat{\sigma}^{\prime})+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{4}(n)~{},$
@@
$R_{\alpha}(\widehat{\sigma}_{J})=\mathbb{E}\left[\sum_{i\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right]+\mathbb{E}\left[\sum_{i\not\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right]~{}.$
@@
$\leq\mathbb{E}\left[D\right]\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+\mathbb{E}\left[\sum_{i\not\in\{1\to c\}}\frac{\left|\widehat{\sigma}^{\prime}(i)-i\right|}{i^{\alpha}}\right]~{}.$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right]\leq\mathbb{E}\left[\sum_{i\in\{1\to c\}}i\right]+\mathbb{E}\left[\sum_{i\in\{1\to c\}}\widehat{\sigma}_{J}(i)\right]~{}.$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}i\right]\leq\mathbb{E}[cD]~{}.$
@@
$\mathbb{E}\left[cD\right]\leq\mathbb{E}\left[c^{2}\right]\leq 2K^{2}~{}.$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}\widehat{\sigma}_{J}(i)\right]\leq\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]~{}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\left(\mathbb{E}\left[\widehat{\sigma}_{J}(1)^{1+\gamma}\right]\right)^{\frac{1}{1+\gamma}}~{}.$ (9)
@@
$\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\leq C\frac{1+\gamma}{\gamma}~{},$ (10)
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq f(\epsilon)\right\}\leq 5\epsilon~{},$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k\right\}\leq 25\frac{\log^{2}(k)}{k}~{},\quad\text{for all}\ k\geq 2,$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)^{1+\gamma}\geq k\right\}=\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k^{\frac{1}{1+\gamma}}\right\}\leq 25\frac{\frac{1}{(1+\gamma)^{2}}\log^{2}(k)}{k^{\frac{1}{1+\gamma}}}~{}.$
@@
$=1+2^{1+\gamma}+25\frac{(1+\gamma)\log^{2}(n)}{\gamma}n^{\gamma}~{}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C\frac{1+\gamma}{\gamma}\left(1+2^{1+\gamma}+25{1+\gamma\over\gamma}\log^{2}(n)n^{\gamma}\right)~{}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C^{\prime}\log^{4}(n)~{}.$

### v3

@@ 1 Introduction
$\widehat{\sigma}(T,\sigma)\overset{\mathcal{L}}{=}\widehat{\sigma}(T^{\sigma^{\prime}},\sigma\circ\sigma^{\prime}),$ (1)
@@
$R_{\alpha}(\widehat{\sigma})\ \stackrel{{\scriptstyle\rm def}}{{=}}\ \mathbb{E}\left[\sum_{i=1}^{n}\frac{|\widehat{\sigma}(i)-\sigma(i)|}{\sigma(i)^{\alpha}}\right],$ (2)
@@
While the risk defined in (2) can be computed for any value $\alpha>0$, we restrict $\alpha$ to a range of values which are relevant for our ordering problem. Specifically, we only consider $\alpha\geq 1$, since for $\alpha<1$ the problem becomes trivial, since even a random permutation has a risk which is minimax optimal up to constant factor (see Appendix A.1). In Theorems 1 and 8, we provide minimax lower bounds for the risk $R_{\alpha}(\widehat{\sigma})$ in the urrt and pa model, for any label-invariant estimator $\widehat{\sigma}$. Then, in Theorems 4 and 9 upper bounds for the risk of the Jordan ordering are obtained. Finally, in Corollaries 5 and 10 we prove that our proposed estimator is minimax optimal up to constant factors, in a non-trivial range of parameters $\alpha$. In the following table, we summarise our findings. For $\alpha\geq 1$, we denote by $R^{*}_{\alpha}$ the optimal risk, and $R_{\alpha}(\widehat{\sigma}_{J})$ the risk of the Jordan ordering.
@@
In the rest of this section, we review previous works and introduce some notation. Then, in Section 2, we analyze the Jordan ordering in the urrt model. Next, we consider the pa model and prove analogous results in Section 4. Finally, in the Supplementary Material, we report the results of our simulation study and compare the empirical performance of the Jordan estimator to alternative methods based on the degree centrality, a peeling method (Navlakha and Kingsford [25, Section 2.3]) and a spectral method commonly used in seriation problems (Recanati et al. [27]).
@@ 1.1 Related work
The vertex arrival-time estimation problem bears some similarity to the seriation problem, though in the former, the dependence is intrinsically related to the tree structure. For example, in a random geometric graph (Gilbert [15]), the seriation problem is to estimate the position of the random points. Since there is no time structure in seriation, different metrics for the error are used, such as the maximum distance between the true and estimated latent position. Examples of methods are provided by Giraud et al. [16]. Another widely studied seriation method consists in ordering latent points by a spectral method on the graph Laplacian (see Supplementary material and Recanati et al. [27] for details). They give guarantees for the quality of their method when the observed adjacency matrix is a perturbed Robinson matrix. The expected adjacency matrices of both urrt and pa trees are Robinson, and therefore the models studied here can be viewed as perturbed Robinson matrices. Nonetheless, none of the above-mentioned papers gives any insight about seriation in urrt and pa trees.
@@ 1.2 Notation
Recall that for a tree $T$ and a vertex $u\in T$, we denote by $(T,u)$ the tree rooted at $u$. For a rooted tree $(T,u)$ and a vertex $v$ we denote by $(T,u)_{v}$ the subtree of $T$ consisting of all vertices $w$ such that $v$ lies on the path connecting $w$ to $u$ (see Figure 1). For two vertices $u,v\in T$, $u\sim v$ means that $u$ is a neighbor of $v$ in $T$ (and reciprocally). In a rooted tree $(T,u)$, we say that $w$ is a child of $v$ if $w$ is in $(T,u)_{v}$. We denote by ${\rm de}_{n}(u)=|(\mathcal{T}_{n},1)_{u}|-1$ the number of descendants of $u$ in $\mathcal{T}_{n}$. For simplicity, we drop the subscript $n$ and use ${\rm de}(u)$ when the size of the tree is fixed and clear from the context.
@@ 2.1 A lower bound
$R^{*}_{\alpha}:=\min_{\widehat{\sigma}\in\Pi_{n}}R_{\alpha}(\widehat{\sigma}),$
@@
In the urrt model, we have, for all $\alpha>0$ and $n\geq 200$,
$R^{*}_{\alpha}\geq\frac{n^{2-\alpha}}{70}.$
@@
$R_{\alpha}(\widehat{\sigma})=\mathbb{E}\left[\sum_{j=1}^{n}\frac{\left|\widehat{\sigma}(j)-\sigma(j)\right|}{\sigma(j)^{\alpha}}\right]=\mathbb{E}\left[\sum_{j=1}^{n}\frac{\left|\widehat{\sigma}\circ\tau(j)-j\right|}{j^{\alpha}}\right],$
@@
$\geq\frac{1}{n^{\alpha}}\sum_{j=\left\lfloor n/2\right\rfloor+1}^{\left\lfloor 3n/4\right\rfloor}\mathbb{E}\left[\big{|}\widehat{\sigma}\circ\tau(j)-j\big{|}+\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\ \left\lfloor\frac{n}{4}\right\rfloor-j\bigg{|}\right].$ (6)
@@
$\Omega_{j}:=\left\{\tau(j)\text{ and }\tau(\left\lfloor n/4\right\rfloor+j)\text{ are leaves, connected to vertices of rank }\leq n/2\right\}.$
@@
    $\prod_{k=j+1}^{\left\lfloor n/4\right\rfloor+j-1}\frac{k-2}{k-1}.$
@@
    $\prod_{k=\left\lfloor n/4\right\rfloor+j+1}^{n}\frac{k-3}{k-1}.$
@@
$\mathbb{P}\left\{\Omega_{j}\right\}$ $=\frac{\left\lfloor n/2\right\rfloor}{j-1}\cdot\frac{\left\lfloor n/2\right\rfloor}{\left\lfloor n/4\right\rfloor+j-1}\cdot\prod_{k=j+1}^{\left\lfloor n/4\right\rfloor+j-1}\frac{k-2}{k-1}\cdot\prod_{k=\left\lfloor n/4\right\rfloor+j+1}^{n}\frac{k-3}{k-1}$
$=\frac{\left\lfloor n/2\right\rfloor}{j-1}\cdot\frac{\left\lfloor n/2\right\rfloor}{\left\lfloor n/4\right\rfloor+j-1}\cdot\frac{j-1}{\left\lfloor n/4\right\rfloor+j-2}\cdot\frac{(\left\lfloor n/4\right\rfloor+j-2)(\left\lfloor n/4\right\rfloor+j-1)}{(n-2)(n-1)},$
@@
$\mathbb{P}\left\{\Omega_{j}\right\}\geq\frac{1}{4}\left(1-\frac{1}{n-1}\right).$ (7)
@@
$\geq$ $\frac{1}{4}\left(1-\frac{1}{n-1}\right)\mathbb{E}\left[\left.\big{|}\widehat{\sigma}\circ\tau(j)-j\big{|}+\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)\bigg{|}\ \right|\ \Omega_{j}\right].$
@@
$\ \ \ \ +\sum_{t\in\mathbb{T}}\mathbb{P}\left\{T=t^{\gamma}\mid\Omega_{j}\right\}\mathbb{E}\left[\left.\bigg{|}\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)\bigg{|}\ \right|\ \Omega_{j},\ T=t^{\gamma}\right],$
@@
$\mathbb{P}\left\{T=t\mid\Omega_{j}\right\}=\mathbb{P}\left\{T=t^{\gamma}\mid\Omega_{j}\right\}.$
@@
$\widehat{\sigma}[T^{\gamma}]\circ\gamma\overset{\mathcal{L}}{=}\widehat{\sigma}[T],$
@@
$\left(\widehat{\sigma}\circ\tau(j)\ \mid\Omega_{j},T=t\right)\overset{\mathcal{L}}{=}\left(\widehat{\sigma}\circ\tau\left(\left\lfloor n/4\right\rfloor+j\right)\ \mid\Omega_{j},T=t^{\gamma}\right),$
@@
$\mathbb{E}\left[|\widehat{\sigma}\circ\tau(j)-j|\ \mid\ \Omega_{j},\ T=t\right]+\mathbb{E}\left[\left.\left|\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left\lfloor\frac{n}{4}\right\rfloor-j\right|\ \right|\ \Omega_{j},\ T=t^{\gamma}\right]\geq\lfloor n/4\rfloor.$
@@
$\mathbb{E}\left[|\widehat{\sigma}\circ\tau(j)-j|+\left|\widehat{\sigma}\circ\tau\left(\left\lfloor\frac{n}{4}\right\rfloor+j\right)-\left\lfloor\frac{n}{4}\right\rfloor-j\right|\right]\geq\frac{n-4}{16}\left(1-\frac{1}{n-1}\right).$
@@
$R_{\alpha}(\widehat{\sigma})\geq\frac{n^{2-\alpha}}{70},$
@@ 2.2 An auxiliary “descendant-ordering” procedure
    $\widehat{\sigma}_{J}(1)\leq 2.5\frac{\log(1/\epsilon)}{\epsilon}.$
@@
$R_{\alpha}(\widehat{\sigma}_{J})\leq R_{\alpha}(\widehat{\sigma}^{\prime})+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{4}(n),$
@@
$R_{\alpha}(\widehat{\sigma}_{J})=\mathbb{E}\left[\sum_{i\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right]+\mathbb{E}\left[\sum_{i\not\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right].$
@@
$\leq\mathbb{E}\left[D\right]\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+\mathbb{E}\left[\sum_{i\not\in\{1\to c\}}\frac{\left|\widehat{\sigma}^{\prime}(i)-i\right|}{i^{\alpha}}\right].$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}\frac{\left|\widehat{\sigma}_{J}(i)-i\right|}{i^{\alpha}}\right]\leq\mathbb{E}\left[\sum_{i\in\{1\to c\}}i\right]+\mathbb{E}\left[\sum_{i\in\{1\to c\}}\widehat{\sigma}_{J}(i)\right].$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}i\right]\leq\mathbb{E}[cD].$
@@
$\mathbb{E}\left[cD\right]\leq\mathbb{E}\left[c^{2}\right]\leq 2K^{2}.$
@@
$\mathbb{E}\left[\sum_{i\in\{1\to c\}}\widehat{\sigma}_{J}(i)\right]\leq\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right].$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\left(\mathbb{E}\left[\widehat{\sigma}_{J}(1)^{1+\gamma}\right]\right)^{\frac{1}{1+\gamma}}.$ (9)
@@
$\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\leq C\frac{1+\gamma}{\gamma},$ (10)
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq f(\epsilon)\right\}\leq 5\epsilon,$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k\right\}\leq 25\frac{\log^{2}(k)}{k},\quad\text{for all}\ k\geq 2,$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)^{1+\gamma}\geq k\right\}=\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k^{\frac{1}{1+\gamma}}\right\}\leq 25\frac{\frac{1}{(1+\gamma)^{2}}\log^{2}(k)}{k^{\frac{1}{1+\gamma}}}.$
@@
$=1+2^{1+\gamma}+25\frac{(1+\gamma)\log^{2}(n)}{\gamma}n^{\gamma}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C\frac{1+\gamma}{\gamma}\left(1+2^{1+\gamma}+25{1+\gamma\over\gamma}\log^{2}(n)n^{\gamma}\right).$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C^{\prime}\log^{4}(n).$